OpenStudy (anonymous):

Frieda earns $7 an hour plus she recives a $50 bonus each month she continues to work at the store. If before taxes she earned $917 on her monthly paycheck, how many hours did she work this month? A) Set up the 2 step algebraic equation that represents the problem.

OpenStudy (anonymous):

what the monkey is a "two step" equation?

OpenStudy (anonymous):

before the bonus she made \(917-50\) dollars divide that amount by 7 to see how many hours she worked (hope she gets a real job soon)

OpenStudy (anonymous):

So could it be 7x+50=917?

OpenStudy (triciaal):

If before taxes she earned $917 on her monthly paycheck, 917 - 50 = amount from the hours worked $7/hr * no of hours = $ earned without the bonus (917- 50 )

OpenStudy (anonymous):

Wait... so what's the equation?

OpenStudy (triciaal):

how many hours did she work this month? let no of hours = x then 7 x = (917-50) solve for x